OpenStudy (anonymous):

A clinical trial tests a method designed to increase the probability of conceiving a girl. In the study 520 babies were born, and 286 of them were girls. Use the sample data to construct a 99% confidence interval estimate of the percentage of girls born. Based on the result, does the method appear to be effective? __< P < __

OpenStudy (amistre64):

what do you believe our hypothesis should be?

OpenStudy (anonymous):

By finding the standard deviation and the mean I know for sure

OpenStudy (amistre64):

those are already presented to yo in the problem, the first step is determining what our hypothesises should be so we know what we are testing for

OpenStudy (amistre64):

i may be confusing this with a hypothesis test ... its simple to misread it :) what is out mean?

OpenStudy (amistre64):

and in this case its called a proportion, the trials produced a proportion of girls born.

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

How can I find that on the TI83?

OpenStudy (amistre64):

theres a 1 sample proportion test id have to find it again, its either a stat test under th elist menu or a side menu under 2nd vars

OpenStudy (anonymous):

Ok thank you

OpenStudy (amistre64):

hit the stat button

OpenStudy (amistre64):

arrow over to "tests" arrow down to the 1 propZtest

OpenStudy (amistre64):

its asks for n, p, and confidence level

OpenStudy (amistre64):

hmm, thats still a hypothesis test tho

OpenStudy (amistre64):

our interval itself is: \[\frac xn \pm z\sqrt{\frac{x(n-x)}{n^3}}\] assuming of course we use the trial data

OpenStudy (amistre64):

your: __ < P < ___ is a little odd to me. it looks like its asking for a confidence interval